OpenStudy (moongazer):

find the particular solution satisfying the initial condition indicated. y' = x e^(y-x^2) when x=0, y=0

OpenStudy (moongazer):

I tried using ln, it seems to go nowhere

ganeshie8 (ganeshie8):

separation of variables

ganeshie8 (ganeshie8):

\[\large y' = x e^{y-x^2} \] \[\large \dfrac{dy}{dx} = \dfrac{x e^{y}}{e^{x^2}} \] \[\large \dfrac{1}{e^y}dy = \dfrac{x }{e^{x^2}} dx \]

ganeshie8 (ganeshie8):

integrate ^^

OpenStudy (moongazer):

wow! I didn't thought of doing that. thanks! :)

ganeshie8 (ganeshie8):

yw :)

OpenStudy (moongazer):

@ganeshie8 just a clarification, is the answer 2e^-y + e^(x^2) = 3 ?

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{e^y}dy = \dfrac{x }{e^{x^2}} dx\] \[\large \int e^{-y}dy = \int x e^{-x^2} dx\]

ganeshie8 (ganeshie8):

\[\large -e^{-y}= -\frac{1}{2} e^{-x^2} + C\] \[\large e^{-x^2}-2e^{-y}= C\]

ganeshie8 (ganeshie8):

thats the general solution, right ?

OpenStudy (moongazer):

yes

ganeshie8 (ganeshie8):

plugin (0, 0) you get C = 1 - 2 = -1

ganeshie8 (ganeshie8):

then the particular solution would be \[\large e^{-x^2}-2e^{-y}= -1\]

OpenStudy (moongazer):

oh, I got the sign incorrectly. can you change the equation into y=ln [2/(1+e^-(x^2))] ? because that's the answer on my book

ganeshie8 (ganeshie8):

yes just isolate the y

ganeshie8 (ganeshie8):

\[\large e^{-x^2}-2e^{-y}= -1 \] \[\large e^{-x^2}+1 = 2e^{-y} \]

ganeshie8 (ganeshie8):

\[\large e^{-x^2}-2e^{-y}= -1 \] \[\large \dfrac{e^{-x^2}+1}{2} = e^{-y} \] \[\large \dfrac{2}{e^{-x^2}+1} = e^{y} \]

ganeshie8 (ganeshie8):

take ln both sides

ganeshie8 (ganeshie8):

\[\large \ln\left( \dfrac{2}{e^{-x^2}+1}\right) = \ln e^{y} \]

ganeshie8 (ganeshie8):

\[\large \ln\left( \dfrac{2}{e^{-x^2}+1}\right) = y\ln e \]

ganeshie8 (ganeshie8):

\[\large \ln\left( \dfrac{2}{e^{-x^2}+1}\right) = y \]

OpenStudy (moongazer):

ok, Thanks! I understood it. :)